# Efficient and faithful reconstruction of dynamical attractors using homogeneous differentiators

### v1

@@ 1 Introduction, Literature Overview and Motivation
We thoroughly review the theory of HDs in Section 2 (as well as in Section S1 of the Supplementary Material), where we also discuss their applicability for attractor reconstruction from time series and provide indications for their computational implementation and practical usage. In Section 3, we conduct our numerical investigation on models of increasing complexity, subject to a variety of noise types, both additive and multiplicative. In Section 3.1, we study the benchmark example of the chaotic Lorenz’63 system [50] and we quantify the performance of the HD compared to a commonly employed noise-reducing method, in terms of quality of the attractor reconstruction and required computational effort. In Section 3.2, we consider the Hindmarsh-Rose (HR) model [51] for neural dynamics and use the attractor reconstruction results obtained from applying the HD on measurements of a single variable to study neural computation properties without access to direct measurements of the membrane potential. In Section 3.3, for the Epileptor model [52] (an extension of the HR model for simulation of seizure-like events in groups of neurons), we investigate how applying the HD on measurements of multiple variables improves the quality of attractor reconstruction in a large-scale system. Finally, in Section 4, we consider real measurement data and apply the HD to reconstruct the dynamic attractor of epileptic brain activity from empirical EEG signals; our results support the use of the HD to preprocess noisy signals and reconstructing attractors of highly nonlinear dynamical systems also based on real noisy data. Section 5 provides a final discussion of the results and of possible future directions.
@@ 2.1 The Homogeneous filtering Differentiator (HD) in continuous and discrete time
Consider a scalar measured signal of the form $f(t)=f_{0}(t)+\eta(t)$, $t\in\mathbb{R}_{+}$, where $f_{0}(t)$ is a (clean) base signal that is corrupted by a noise $\eta(t)$. Let us denote by $f^{(i)}_{0}(t)$ the $i$-th time derivative of $f_{0}(t)$. For a pre-defined number $n_{d}\in\mathbb{N}_{0}$, the differentiation problem consists of estimating $f^{(0)}_{0}(t),\dots,f^{(n_{d})}_{0}(t)$ in finite-time from the measured signal $f(t)$. We make the following assumptions:
@@
where $w=(w_{1},\dots,w_{n_{f}})^{\top}$ is the state of the filtration block, $z=(z_{0},\dots,z_{n_{d}})^{\top}$ is the state of the differentiation block, $f$ is the measured (noisy) signal and $\Omega_{n_{d},n_{f}}$, $D_{n_{d},n_{f}}$ are appropriate vector-valued functions (whose complete expression is reported above Eq. (S1) in the Supplementary Material). The scheme is initiated with the parameters $n_{d}$, $n_{f}$ and $L_{0}$ supplied by the user. The additional parameters in vector $\lambda=(\tilde{\lambda}_{0},\dots,\tilde{\lambda}_{n_{d}+n_{f}})$ are extracted from a pre-specified array and can be explicitly provided for $n_{d}+n_{f}\leq 12$ (however, a suitable set of parameters is proven to exist for all possible $n_{d}$ and $n_{f}$, via a recursive construction [26]; see Section S1.1 in the Supplementary Material). Intuitively speaking, the $w$-system (filtration block) is responsible for eliminating the noise components $\eta_{1},\dots,\eta_{n_{f}}$ by filtering them out of the signal $f$, so that in finite-time $w_{1}(t)\approx f_{0}(t)+\eta_{0}(t)$ holds. Then, $w_{1}$ is fed into the $z$-system, which performs the differentiation and produces estimates $z_{i}(t)\approx f_{0}^{(i)}(t),\ i=0,\dots,n_{d}$ in finite-time and without resorting to differentiation of the noise components. An important feature of the HD, crucial since the true nature of the corrupting noise $\eta(t)$ is not known, is that increasing the filtration degree $n_{f}$ does not deteriorate the performance; by taking larger values of $n_{f}$, one can filter out more complex noises and potentially reduce the differentiation error. Note that we are not making any assumptions on the smoothness of the noise $\eta$, as those are not required for the HD to be applied; in fact, homogeneous differentiators provide estimates of the base signal and its time derivatives with guaranteed accuracy regardless of whether the noise is smooth or non-smooth, as no derivatives of $\eta$ are taken in the dynamics of the HD.
@@
The sampled noise has $n_{f}+1\in\mathbb{N}_{0}$ components, $\eta(t_{s})=\sum_{j=0}^{n_{f}}\eta_{j}(t_{s})$, where $\eta_{j}(t_{s})=\underline{\eta}_{j}(t_{s})+\bar{\eta}_{j}(t_{s}),\ j=0,\dots,n_{f}$. The discretely sampled signals $\underline{\eta}_{j}(t_{s})$ are of the global sampling filtering order $j$ with $j$-th order global integral sampling magnitude $\underline{\epsilon}_{j}$ (see Definition 3, Section S1.2 in the Supplementary Material); $\bar{\eta}_{0}=0$ and each noise component $\bar{\eta}_{j}$ is of the (continuous-time) global filtering order $j$ with $j$-th order global integral magnitude $\bar{\epsilon}_{j}$ for $j=1,\dots,n_{f}$ (see Definition 1, Section S1 in the Supplementary Material). Each noise component ${\bar{\eta}}_{j}(t)$ is also absolutely continuous with $\left|\dot{\bar{\eta}}_{j}(t)\right|\leq L_{\eta_{j}}$ for all $j=1,\dots,n_{f}$ for some $L_{\eta_{j}}>0$.
@@ 2.2 Homogeneous-Differentiator-based attractor reconstruction
Differently from the the Schreiber-Grassberger algorithm and its derivations, which require multiple computationally expensive operations for each data point, the HD can be implemented easily and with computational efficiency by a finite-difference scheme. Also, the HD is effective in the presence of any measurable and essentially bounded noise signal $\eta(t)$, regardless of the specific system generating the base signal $\bar{y}(t)$; strong theoretical guarantees and error bounds are obtained when the noise is additive, and hence $y(t)=\bar{y}(t)+\eta(t)$. Moreover, as it can be seen in Figure 1, the HD setup only requires to choose three parameters: the number $n_{d}$ of reconstructed derivatives; a bound $L_{0}$ on the Lipschitz constant of the $n_{d}$-th derivative; the filtration degree $n_{f}$, which can be always increased without deteriorating the accuracy of the estimation [28] (see also the discussion in Section 2.1). An estimate of $L_{0}$ is typically not available a priori; however, to tune the parameter, one can assess the quality of the HD output obtained for increasing values of $L_{0}$, and note that $L_{0}$ values that are too large lead to chattering (i.e., small-scale high-frequency oscillations) in the HD output [28]; see Supplementary Figure S3. To further address possible chattering issues [29], the HD output can be post-processed to filter out undesired high-frequency oscillations; to this aim, here we use the Savitzky-Golay (SG) filter [54], which has zero phase, very flat responses in the passband, and moderate attenuation in the stopband, and thus leaves the main characteristics of the signal undistorted, while high-frequency components are reduced. Post-processing the HD output with the SG filter appears successful in practice, but the theoretical accuracy guarantees that are valid for the HD output do not extend to the SG filter output (because the SG filter does not enjoy the same homogeneity properties that characterise the HD).
@@ 3.1 Lorenz’63 system: chaotic dynamics
For a quantitative comparison between the different methods, we compute the relative errors $E_{R,i}$ in (3). Table 1, top, summarises the most relevant statistics of the distribution of $E_{R,i}(k)$ (for the full distributions of $E_{R,i}(k)$, we refer to the Supplementary Figure S5). The Differentiator significantly outperforms Schreiber-Grassberger noise reduction, which already suppresses the discrepancies associated with direct embedding of noisy measurements, by providing more than tenfold error reduction in the case of Gaussian and unbounded noise and an even stronger reduction in the case of harmonic noise. The performance of the Differentiator further improves after SG post-filtering. The advantage of the Differentiator is even more pronounced in the case of multiplicative noise.
@@ 3.3 Epileptor model: measuring multiple variables to monitor epileptic seizures
Figure 8: The Differentiator enables de-noising and faithful attractor reconstruction for the Epileptor system affected by various types of additive and multiplicative noise, in the state-space coordinates. Reconstructed attractor of the Epileptor model (11) from measurements of the signals $x_{1}$, $x_{4}$ and $x_{3}$ affected by either additive or multiplicative noise, obtained through the HD-based method. The Differentiator parameters are set as $n_{d}=2$, $n_{f}=10$, $L_{0}=50$. Rows, from top to bottom: noise-free signals $x_{1}$, $x_{4}$ and $x_{3}$ to be measured, along with the true attractor of the system (black) in the $\{x_{1},x_{3},x_{4}\}$ sub-space; attractor reconstruction with additive noise; attractor reconstruction with multiplicative noise. Columns, from left to right: both in the case of additive and of multiplicative noise, reconstruction of the attractor (in the same space) from measurements corrupted by multiplicative noise $\eta_{H}(t)$ as in (5) (green); $\eta_{U}(t)$ as in (6) (red); $\eta_{1}(t)\sim\mathcal{N}(0,0.01)$ (blue); $\eta_{2}(t)\sim\mathcal{N}(0,0.1)$ (orange); $\eta_{3}(t)\sim\mathcal{N}(0,1)$ (violet).
@@ 4 Reconstructing Unknown Attractors from Empirical Data
Furthermore, Figure 10 shows the signal de-noising and attractor reconstruction (in three-dimensional time-delay coordinates as above) for the segmented $2,000$ Hz time series of background and seizure activity. These are different events occurred on a single fish, during the recordings (other events for the same fish are shown in the Supplementary Figure S27). As before, the Differentiator suppresses most of the measurement noise, ideally unveiling the core dynamics and highlighting the key topological differences between the two types of events. Background events are characterised by an agglomerate of points that, except for punctual and sparse spikes, are akin to typical attractors of semi-random signals; this is consistent with the neurological hypothesis that the global brain dynamics is overall incoherent [71, 72]. For seizure events, the Differentiator highlights a clear structure that evolves as a semi-loop and suggests some form of pattern or coherence; this is in line with the neurological hypothesis that seizures emerge from spontaneous alignment of neural firing [73, 74]. Moreover, these semi-loop structures may be linked to those loops observed in Figure 9, which are captured by dynamical models such as the Epileptor. If our hypothesis were confirmed, the Differentiator would emerge as the tool of choice to help verify models and topological characteristics of seizure events, thus improving their detection and prediction. The considered dataset does not provide a mapping between the segmented time series with labelling of conditions and the full unsegmented time series; hence, we cannot fully test this interesting hypothesis, which is left for future investigation.
Overall, our preliminary results demonstrate the impressive capability of the Differentiator to process and de-noise empirical signals, and to reconstruct cleaned-up attractors, thus highlighting their key topological features that would otherwise be obscured by unknown noise. In turn, unveiling these features enables the development of new hypothesis, the verification of dynamical models from data, and a deeper subsequent analysis.
@@ Acknowledgments
Code availability and Supplementary Material

### v2

@@ 1 Introduction, Literature Overview and Motivation
We thoroughly review the theory of HDs in Section 2 (as well as in Section S1 of the Supplementary Material), where we also discuss their applicability for attractor reconstruction from time series and provide indications for their computational implementation and practical usage. In Section 3, we conduct our numerical investigation on models of increasing complexity, subject to a variety of noise types, both additive and multiplicative. In Section 3.1, we study the benchmark example of the chaotic Lorenz’63 system [50] and we quantify the performance of the HD compared to a commonly employed noise-reducing method, in terms of quality of the attractor reconstruction and required computational effort. In Section 3.2, we consider the Hindmarsh-Rose (HR) model [51] for neural dynamics and use the attractor reconstruction results obtained from applying the HD on measurements of a single variable to study neural computation properties without access to direct measurements of the membrane potential. In Section 3.3, for the Epileptor model [52] (an extension of the HR model for simulation of seizure-like events in groups of neurons), we investigate how applying the HD on measurements of multiple variables improves the quality of attractor reconstruction in a large-scale system. Finally, in Section 4, we consider real measurement data and apply the HD to reconstruct the dynamic attractor of epileptic brain activity from empirical EEG signals; our results support the use of the HD to preprocess noisy signals and reconstruct attractors of highly nonlinear dynamical systems also based on real noisy data. Section 5 provides a final discussion of the results and of possible future directions.
@@ 2.1 The Homogeneous filtering Differentiator (HD) in continuous and discrete time
Consider a scalar measured signal of the form $f(t)=f_{0}(t)+\eta(t)$, $t\in\mathbb{R}_{+}$, where $f_{0}(t)$ is a (clean) base signal that is corrupted by a noise $\eta(t)$. Let us denote by $f^{(i)}_{0}(t)$ the $i$-th time derivative of $f_{0}(t)$. For a pre-defined number $n_{d}\in\mathbb{N}_{0}$, the differentiation problem consists of estimating $f^{(0)}_{0}(t),\dots,f^{(n_{d})}_{0}(t)$ in finite time from the measured signal $f(t)$. We make the following assumptions:
@@
where $w=(w_{1},\dots,w_{n_{f}})^{\top}$ is the state of the filtration block, $z=(z_{0},\dots,z_{n_{d}})^{\top}$ is the state of the differentiation block, $f$ is the measured (noisy) signal and $\Omega_{n_{d},n_{f}}$, $D_{n_{d},n_{f}}$ are appropriate vector-valued functions (whose complete expression is reported above Eq. (S1) in the Supplementary Material). The scheme is initiated with the parameters $n_{d}$, $n_{f}$ and $L_{0}$ supplied by the user. The additional parameters in vector $\lambda=(\tilde{\lambda}_{0},\dots,\tilde{\lambda}_{n_{d}+n_{f}})$ are extracted from a pre-specified array and can be explicitly provided for $n_{d}+n_{f}\leq 12$ (however, a suitable set of parameters is proven to exist for all possible $n_{d}$ and $n_{f}$, via a recursive construction [26]; see Section S1.1 in the Supplementary Material). Intuitively speaking, the $w$-system (filtration block) is responsible for eliminating the noise components $\eta_{1},\dots,\eta_{n_{f}}$ by filtering them out of the signal $f$, so that in finite time $w_{1}(t)\approx f_{0}(t)+\eta_{0}(t)$ holds. Then, $w_{1}$ is fed into the $z$-system, which performs the differentiation and produces estimates $z_{i}(t)\approx f_{0}^{(i)}(t),\ i=0,\dots,n_{d}$ in finite time and without resorting to differentiation of the noise components. An important feature of the HD, crucial since the true nature of the corrupting noise $\eta(t)$ is not known, is that increasing the filtration degree $n_{f}$ does not deteriorate the performance; by taking larger values of $n_{f}$, one can filter out more complex noises and potentially reduce the differentiation error. Note that we are not making any assumptions on the smoothness of the noise $\eta$, as those are not required for the HD to be applied; in fact, homogeneous differentiators provide estimates of the base signal and its time derivatives with guaranteed accuracy regardless of whether the noise is smooth or non-smooth, as no derivatives of $\eta$ are taken in the dynamics of the HD.
@@
The sampled noise has $n_{f}+1\in\mathbb{N}$ components, $\eta(t_{s})=\sum_{j=0}^{n_{f}}\eta_{j}(t_{s})$, where $\eta_{j}(t_{s})=\underline{\eta}_{j}(t_{s})+\bar{\eta}_{j}(t_{s}),\ j=0,\dots,n_{f}$. The discretely sampled signals $\underline{\eta}_{j}(t_{s})$ are of the global sampling filtering order $j$ with $j$-th order global integral sampling magnitude $\underline{\epsilon}_{j}$ (see Definition 3, Section S1.2 in the Supplementary Material); $\bar{\eta}_{0}=0$ and each noise component $\bar{\eta}_{j}$ is of the (continuous-time) global filtering order $j$ with $j$-th order global integral magnitude $\bar{\epsilon}_{j}$ for $j=0,\dots,n_{f}$ (see Definition 1, Section S1 in the Supplementary Material). Each noise component ${\bar{\eta}}_{j}(t)$ is also absolutely continuous with $\left|\dot{\bar{\eta}}_{j}(t)\right|\leq L_{\eta_{j}}$ for all $j=0,\dots,n_{f}$ for some $L_{\eta_{j}}>0$.
@@ 2.2 Homogeneous-Differentiator-based attractor reconstruction
Differently from the the Schreiber-Grassberger algorithm and its derivations, which require multiple computationally expensive operations for each data point, the HD can be implemented easily and with computational efficiency by a finite-difference scheme. Also, the HD is effective in the presence of any measurable and essentially bounded noise signal $\eta(t)$, regardless of the specific system generating the base signal $\bar{y}(t)$; strong theoretical guarantees and error bounds are obtained when the noise is additive, and hence $y(t)=\bar{y}(t)+\eta(t)$. Moreover, as it can be seen in Figure 1, the HD setup only requires to choose three parameters: the number $n_{d}$ of reconstructed derivatives; a bound $L_{0}$ on the Lipschitz constant of the $n_{d}$-th derivative; the filtration degree $n_{f}$, which can be always increased without deteriorating the accuracy of the estimation [28] (see also the discussion in Section 2.1). An estimate of $L_{0}$ is typically not available a priori; however, to tune the parameter, one can assess the quality of the HD output obtained for increasing values of $L_{0}$, and note that $L_{0}$ values that are too large lead to chattering (i.e., small-scale high-frequency oscillations) in the HD output [28]; see Supplementary Figure S3. To further address possible chattering issues [29], the HD output can be post-processed to filter out undesired high-frequency oscillations; to this aim, here we use the Savitzky-Golay (SG) filter [54], which has zero phase, very flat response in the passband, and moderate attenuation in the stopband, and thus leaves the main characteristics of the signal undistorted, while high-frequency components are reduced. Post-processing the HD output with the SG filter appears successful in practice, but the theoretical accuracy guarantees that are valid for the HD output do not extend to the SG filter output (because the SG filter does not enjoy the same homogeneity properties that characterise the HD).
@@ 3.1 Lorenz’63 system: chaotic dynamics
For a quantitative comparison between the different methods, we compute the relative errors $E_{R,i}$ as in (3). Table 1, top, summarises the most relevant statistics of the distribution of $E_{R,i}(k)$ (for the full distributions of $E_{R,i}(k)$, we refer to the Supplementary Figure S5). The Differentiator significantly outperforms Schreiber-Grassberger noise reduction, which already suppresses the discrepancies associated with direct embedding of noisy measurements, by providing more than tenfold error reduction in the case of Gaussian and unbounded noise and an even stronger reduction in the case of harmonic noise. The performance of the Differentiator further improves after SG post-filtering. The advantage of the Differentiator is even more pronounced in the case of multiplicative noise.
@@ 3.3 Epileptor model: measuring multiple variables to monitor epileptic seizures
Figure 8: The Differentiator enables de-noising and faithful attractor reconstruction for the Epileptor system affected by various types of additive and multiplicative noise, in the state-space coordinates. Reconstructed attractor of the Epileptor model (11) from measurements of the signals $x_{1}$, $x_{4}$ and $x_{3}$ affected by either additive or multiplicative noise, obtained through the HD-based method. The Differentiator parameters are set as $n_{d}=2$, $n_{f}=10$, $L_{0}=50$. Rows, from top to bottom: noise-free signals $x_{1}$, $x_{4}$ and $x_{3}$ to be measured, along with the true attractor of the system (black) in the $\{x_{1},x_{3},x_{4}\}$ sub-space; attractor reconstruction with additive noise; attractor reconstruction with multiplicative noise. Columns, from left to right: both in the case of additive and of multiplicative noise, reconstruction of the attractor (in the same space) from measurements corrupted by noise $\eta_{H}(t)$ as in (5) (green); $\eta_{U}(t)$ as in (6) (red); $\eta_{1}(t)\sim\mathcal{N}(0,0.01)$ (blue); $\eta_{2}(t)\sim\mathcal{N}(0,0.1)$ (orange); $\eta_{3}(t)\sim\mathcal{N}(0,1)$ (violet).
@@ 4 Reconstructing Unknown Attractors from Empirical Data
Furthermore, Figure 10 shows the signal de-noising and attractor reconstruction (in three-dimensional time-delay coordinates as above) for the segmented $2,000$ Hz time series of background and seizure activity. These are different events occurred on a single fish, during the recordings (other events for the same fish are shown in the Supplementary Figure S27). As before, the Differentiator suppresses most of the measurement noise, ideally unveiling the core dynamics and highlighting the key topological differences between the two types of events. Background events are characterised by an agglomerate of points that, except for punctual and sparse spikes, are akin to typical attractors of semi-random signals; this is consistent with the neurological hypothesis that the global brain dynamics is overall incoherent [71, 72]. For seizure events, the Differentiator highlights a clear structure that evolves as a semi-loop and suggests some form of pattern or coherence; this is in line with the neurological hypothesis that seizures emerge from spontaneous alignment of neural firing [73, 74]. Moreover, these semi-loop structures may be linked to the loops observed in Figure 9, which are captured by dynamical models such as the Epileptor. If our hypothesis were confirmed, the Differentiator would emerge as the tool of choice to help verify models and topological characteristics of seizure events, thus improving their detection and prediction. The considered dataset does not provide a mapping between the segmented time series with labelling of conditions and the full unsegmented time series; hence, we cannot fully test this interesting hypothesis, which is left for future investigation.
Overall, our preliminary results demonstrate the impressive capability of the Differentiator to process and de-noise empirical signals, and to reconstruct cleaned-up attractors, thus highlighting their key topological features that would otherwise be obscured by unknown noise. In turn, unveiling these features enables the development of new hypotheses, the verification of dynamical models from data, and a deeper subsequent analysis.
@@ Acknowledgments
Code availability